8. A shot-putter throws the shot with an initial speed of 15.5 m/s at 34.0° angle to the
horizontal. Calculate the horizontal distance travelled by the shot if it leaves the
athlete’s hand at height of 2.20m above the ground. Determine the time flight the shot
stayed in air.
1
Expert's answer
2020-05-11T20:11:51-0400
"y=h+v\sin{30}t-gt^2=0\\ 2.2+15.5\sin{34}t-9.8t^2=0"

"t=1.09\ s"


"x=vt\cos{34}=(15.5)(1.09)\cos{34}=14\ m"